# Multiscale Hodge Scattering Networks for Data Analysis

### v5

@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). For example, in the COLLAB dataset [52], each graph represents a network of coauthors of a scientific paper. Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in [4, 25], which were based on geometric scattering methods.
We compute permutation-invariant input features based only on topological information obtainable from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. Following the methodology developed in [4], we set $(J,M,Q)=(4,2,4)$. For the node signals, we first compute the eccentricity and clustering coefficient [HARRIS-ETAL, Sec. 1.2] of each node. For each node signal, the number of parameters (MHST coefficients) are 64 via the formula $Q\sum_{k=0}^{M}\binom{J+1}{k}$, hence 128 parameters after concatenating them. For the edge signals, we use the average of the eccentricities of the head and tail nodes of each edge and the number of non-zero off-diagonal terms in the combinatorial Hodge-Laplacian (each such term corresponds to a $1$-adjacent edge [29, Sec. 4.1]).
@@
We compare these results with those obtained by the geometric scattering network (with Diffusion Wavelets) using SVM (GS-SVM) as well as several popular GNN models including the graph convolution network (GCN) [40], universal graph transform (UGT) [41], dynamic graph CNN (DGCNN) [42], graph attention network (GAT) [43], and graph feature network (GFN) [44]. For each competing method, we reproduce the results in the significant figures reported in their original publications; we report to 2 decimal places for our method. More information on the benchmark datasets can be found in A. We remark that, as of this article’s writing, this collection of networks achieves state-of-the-art results on these datasets according to the Papers with Code Leaderboards [45]. Further details on these datasets and their associated classification problems are presented in  A and the references therein.
Graph Node Scattering Edge Scattering Combo GS-SVM GCN UGT DGCNN GAT GFN
COLLAB 70.84 78.34 80.39 79.94 79.00 77.84 73.76 75.80 81.50
DD 60.67 68.73 72.71 - - 80.23 79.37 - 79.37
IMDB-B 72.70 70.60 73.10 71.20 74.00 77.04 70.03 70.50 73.40
IMDB-M 44.40 47.13 49.68 48.73 51.90 53.60 47.83 47.8 51.80
MUTAG 85.78 86.31 85.78 83.50 85.60 80.23 79.37 89.40 85.83
PROTEINS 73.57 73.04 75.35 74.11 76.00 78.53 75.54 74.70 76.46
PTC 62.85 67.71 68.28 63.94 64.20 69.63 58.59 66.70 66.60
@@
Although our MHSNs do not achieve state-of-the-art results on these datasets, they are very competitive with only a small fraction of the learnable parameters. Moreover, the number of learnable parameters in our models is not tied to the graph size and depends only on the order of the scattering $M$ and the number of moments $Q$ computed. For example, Table 4 compares our methods with the UGT and the GFN, which are the state-of-the-art methods for various graph classification problems. These methods each require more than half a million parameters for some cases (867K for UGT) to achieve results similar to ours, requiring only 256 parameters to learn. As a result, our MHSNs can be implemented and trained on a consumer-level laptop, whereas many of these competing GNNs require specialized hardware.
Hodge Scattering + SVM UGT GFN
@@
Table 4: Comparison of MHSN and state of the art graph classification networks in accuracy and number of learnable parameters
@@ 7 Molecular Dynamics
We extract a rich set of features for each structure (i.e., a pose or conformation of a molecule) using our MHSNs (without pooling) and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate and predict the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a $k$NN-graph (with $k=5$) as a template simplicial complex. Note that by using this simplicial complex, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in [47, 48]. Unlike the domain classification problems in Section 6, the geometry of the simplicial complex is fixed, so rather than use its geometrically-invariant descriptors, we need to begin with signals that encode the position information of molecules. We then generate the node and edge signals of both training and test sets as follows. First we compute the Euclidean distance matrix of (i.e the Gram matrix of the point-cloud, measured in the Euclidean distance) the node coordinates of each snapshot and assign the corresponding column vector of the distance matrix as its node features. This generates a number-of-atoms-dimensional node signal for each node, for each snapshot. For an edge signal, we extract edge lengths from the above distance matrix and create a diagonal matrix of edge lengths. Then, we assign the corresponding column vector of this diagonal matrix as its edge features. As with the node-based signal, this gives us number-of-edges-dimensional signal to input into our MHSN. For this experiment we use $(J,M,Q)=(4,2,3)$. We do not use simplices of dimension $\kappa>1$ and not set $J>4$ and $Q>3$ because the molecules are too small.
Table 5 shows our results for aspirin (21 atoms) and paracetamol (20 atoms) molecules. We compare our MHSNs with several state-of-the-art GNN approaches designed specifically for processing molecular dynamics, including SchNet [47], PaiNN [49], and two variants of Special Orthogonal Networks (SO3Nets) [50, 48]. We report both the mean absolute error (MAE) and root mean square error (RMSE) of the energy prediction, which are the standard metrics in the computational chemistry literature. Our MHSNs perform competitively with these approaches while employing roughly 1% as many learnable parameters as the competing methods. Additionally, we again observe that our edge-based analysis outperforms the node-based analysis. This demonstrates that higher-dimensional simplex analysis can be more powerful than node-only approaches, even in cases where the underlying graph may not have many higher-dimensional structures.
Diffusion+SVR HGLET+SVR GHWT+SVR SchNet PaiNN SO3Net I SO3Net II
Feature Type Node Edge Both Node Edge Both Node Edge Both
@@
MAE 4.856 3.132 3.267 4.884 3.135 3.285 4.928 3.075 3.225 13.5 3.8 3.8 2.6
RMSE 6.181 4.144 4.314 6.215 4.129 4.407 6.213 4.123 4.316 18.3 5.9 5.7 3.8
# Parameters 924 3784 4708 924 3784 4708 924 3784 4708 $\sim$ 432k $\sim$ 341k $\sim$ 283k $\sim$ 341k
@@
MAE 4.609 2.715 2.795 4.723 2.643 2.710 4.748 2.624 2.699 8.4 2.1 2.2 1.4
RMSE 5.860 3.418 4.116 5.964 3.338 3.424 5.961 3.299 3.408 11.2 2.9 3.0 1.9
# Parameters 924 3784 $4$ 924 3784 $4$ 924 3784 $4$ $\sim$432k $\sim$ 341k $\sim$283k $\sim$341k
Table 5: Comparison of the performance of our MHSNs and the other state-of-the-art GNNs for potential energy prediction. We report the accuracy via MAE and RMSE as well as the number of trainable parameters in each network.
@@ 8 Conclusion
In this article, we proposed the Multiscale Hodge Scattering Transforms/Networks (MHSTs/MHSNs) for robust feature extraction from signals on simplicial complexes that can be used in classification and regression problems, fully utilizing our multiscale basis dictionaries on such simplicial complexes, i.e., $\kappa$-HGLET and $\kappa$-GHWT dictionaries. Our MHSTs/MHSNs also have pooling options for the scattering transform outputs: no-pooling; local-pooling; and global-pooling. Such options allow our tools to apply for various classification and regression problems on simplicial complexes ranging from classification of signals recorded on simplicial complexes to classification of type of simplicial complexes (i.e., domain classification problems) to regression of potential energies in molecular dynamics. We demonstrated that MHSNs provide comparable results with those by the state-of-the-art GNNs with up to a two-order of magnitude reduction in number of learnable parameters. We strongly believe that our success here comes from the structure and organization of our multiscale basis dictionaries that are conveniently arranged in terms of scales and locations, which are suitable and quite amenable for generating scattering transform coefficients.
We plan to investigate how we can interpret the MHST coefficients that are identified as important by classification methods such as the LRMs. Because of the nonlinearities used in the MHSTs, converting the MHST coefficients to the features in the primal/original domain is difficult. Along this direction, we plan to examine the optimization method proposed by Weber [51, Chap. 4], which synthesizes an input signal that generates the significant scattering transform coefficients at the specified coefficient indices. In a related line of research, we will explore how the MHST coefficients can be used to identify important relationships within graphs that can be used to narrow the training space of various attention mechanisms/graph transformers for large-scale problems.
@@ Acknowledgments
This research was partially supported by the US National Science Foundation grants DMS-1912747 and CCF-1934568 as well as the US Office of Naval Research grant N00014-20-1-2381. N. S. also thanks Stéphane Mallat and Gabriel Peyré for their hospitality and support during his sabbatical stay (Oct. 2021–Jan. 2022) at the Centre Sciences des Données, Ecole Normale Supérieure (Ulm).
Appendix A Description of Domain Classification Datasets

### v6

@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). For example, in the COLLAB dataset [40, 41], each graph represents a network of coauthors of a scientific paper. Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in [5, 28], which were based on geometric scattering methods.
We compute permutation-invariant input features based only on topological information obtainable from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. Following the methodology developed in [5], we set $(J,M,Q)=(4,2,4)$. For the node signals, we first compute the eccentricity and clustering coefficient [42, Sec. 1.2] of each node. For each node signal, the number of parameters (MHST coefficients) are 64 via the formula $Q\sum_{k=0}^{M}\binom{J+1}{k}$, hence 128 parameters after concatenating them. For the edge signals, we use the average of the eccentricities of the head and tail nodes of each edge and the number of non-zero off-diagonal terms in the combinatorial Hodge-Laplacian (each such term corresponds to a $1$-adjacent edge [31, Sec. 4.1]).
@@
We compare these results with those obtained by the geometric scattering network (with Diffusion Wavelets) using SVM (GS-SVM) as well as several popular GNN models including the graph convolution network (GCN) [43], universal graph transform (UGT) [44], dynamic graph CNN (DGCNN) [45], graph attention network (GAT) [46], and graph feature network (GFN) [47]. For each competing method, we reproduce the results in the significant figures reported in their original publications; we report to 2 decimal places for our methods. More information on the benchmark datasets can be found in C. We remark that, as of this article’s writing, this collection of networks achieves SotA results on these datasets according to the Papers with Code Leaderboards [48]. Further details on these datasets and their associated classification problems are presented in C and the references therein.
HGLET+SVM GHWT+SVM GS-SVM GCN UGT DGCNN GAT GFN
Graph Node Edge Combo Node Edge Combo
COLLAB 70.84 78.34 80.39 68.94 78.34 78.34 79.94 79.00 77.84 73.76 75.80 81.50
$D$ 60.67 68.73 72.71 72.20 65.76 69.32 - - 80.23 79.37 - 79.37
IMDB-B 72.70 70.60 73.10 62.50 61.60 61.90 71.20 74.00 77.04 70.03 70.50 73.40
IMDB-M 44.40 47.13 49.68 44.33 39.07 39.73 48.73 51.90 53.60 47.83 47.80 51.80
MUTAG 85.78 86.31 85.78 81.58 76.84 79.95 83.50 85.60 80.23 79.37 89.40 85.83
PROTEINS 73.57 73.04 75.35 71.52 66.70 71.61 74.11 76.00 78.53 75.54 74.70 76.46
PTC 62.85 67.71 68.28 52.29 55.71 51.71 63.94 64.20 69.63 58.59 66.70 66.60
@@
Although our MHSNs do not achieve SotA results on these datasets, they are very competitive with only a small fraction of the learnable parameters. Moreover, the number of learnable parameters in our models is not tied to the graph size and depends only on the order of the scattering $M$ and the number of moments $Q$ computed. For example, Table 4 compares our $\kappa$-HGLET-based MHSN with the UGT and the GFN, which are the SotA methods for various graph classification problems. These methods each require more than half a million parameters for some cases (867K for UGT) to achieve results similar to ours, requiring only 256 parameters to learn. As a result, our MHSNs can be implemented and trained on a consumer-level laptop, whereas many of these competing GNNs require specialized hardware.
$\kappa$-HGLET + SVM UGT GFN
@@
Table 4: Comparison of MHSN and the SotA graph classification networks in accuracy and number of learnable parameters
@@ 7 Molecular Dynamics
We extract a rich set of features for each structure (i.e., a pose or conformation of a molecule) using our MHSNs (without pooling) and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate and predict the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a kkNN-graph (with $k=5$) as a template simplicial complex. Note that by using this simplicial complex, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in [50, 51]. Unlike the domain classification problems in Section 6, the geometry of the simplicial complex is fixed, so rather than using its geometrically-invariant descriptors, we need to begin with signals that encode the position information of molecules for each $\kappa$ which we wish to use in the MSHN.
First we compute the Euclidean distance matrix of (i.e., the Gram matrix of the point-cloud, measured in the Euclidean distance) the node coordinates of each snapshot and assign the corresponding column vector of the distance matrix as its node features. This generates a number-of-atoms-dimensional node signal for each node, for each snapshot.
For an edge signal, we extract edge lengths from the above distance matrix and create a diagonal matrix of edge lengths. Then, we assign the corresponding column vector of this diagonal matrix as its edge features. As with the node-based signal, this gives us number-of-edges-dimensional signal to input into our MHSN. We repeat this process using area and volume of the $\kappa=2$ and $\kappa=3$-simplices for the corresponding input singals of $\kappa$-simplices.
We use the setting $(J,M,Q)=(4,2,3)$ to generate the scattering features and use SVR on our final layer output. Finally, we tune the hyperparameters in using the same cross validation scheme and settings in 6, which was orignally proposed in [5].
Table 5 shows our results for aspirin (21 atoms) and paracetamol (20 atoms) molecules. We compare our MHSNs with several SotA GNN approaches designed specifically for processing molecular dynamics, including SchNet [50], PaiNN [52], and two variants of Special Orthogonal Networks (SO3Nets) [53, 51]. We report both the mean absolute error (MAE) and root mean square error (RMSE) of the energy prediction, which are the standard metrics in the computational chemistry literature. Our MHSNs perform competitively with these approaches, while employing roughly 1% as many learnable parameters as competing methods. Additionally, we observe the edge and triangle-based analyses outperform the node-based analysis. This demonstrates that higher-dimensional simplex analysis can be more powerful than node-only approaches, even in cases where the underlying molecular graph may not have many higher-dimensional structures. We also observe that a concatenation of these features, which we refer to in Table 5 as ‘Combo’, outperforms any single $\kappa$-based analysis, just as we observed in the experiments in Section 6.
@@
Diff+SVR HGLET+SVR GHWT+SVR SchNet PaiNN SO3Net I SO3Net II
Feat Node Edge Triangle Pyramid Combo Node Edge Triangle Pyramid Combo Node Edge Triangle Pyramid Combo
MAE 4.856 3.132 2.93 3.857 2.468 4.884 3.135 3.016 3.876 2.651 4.928 3.075 3.01 3.842 2.804 13.5 3.8 3.8 2.6
RMSE 6.181 4.144 3.784 4.937 3.226 6.215 4.129 3.905 4.976 3.473 6.213 4.123 3.883 4.964 3.613 18.3 5.9 5.7 3.8
# Paramters 924 3784 6512 5764 16984 924 3784 6512 5764 16984 924 6512 6512 5764 16984 432k 341k 283k 341k
@@
Diff+SVR HGLET+SVR GHWT+SVR SchNet PaiNN SO3Net I SO3Net II
Feat Node Edge Triangle Pyramid Combo Node Edge Triangle Pyramid Combo Node Edge Trianlge Pyramid Combo
MAE 4.609 2.715 2.77 3.947 2.235 4.723 2.643 2.896 3.995 2.302 4.748 2.624 2.843 3.928 2.309 8.4 2.1 2.2 1.4
RMSE 5.86 3.418 3.596 4.961 2.83 5.964 3.338 3.709 4.99 2.877 5.961 3.299 3.679 4.916 2.932 11.2 2.9 3 1.9
# Paramters 880 3564 5808 4752 15004 880 3564 5808 4752 15004 880 3564 5808 4752 15004 432k 341k 283k 341k
Table 5: Comparison of the performance of our MHSNs and the other SotA GNNs for potential energy prediction. We report the accuracy via MAE and RMSE as well as the number of trainable parameters in each network.
@@ 8 Conclusion
In this article, we proposed the Multiscale Hodge Scattering Transforms/Networks (MHSTs/MHSNs) for robust feature extraction from signals on simplicial complexes, applicable to both classification and regression tasks. Our approach fully leverages our multiscale basis dictionaries on simplicial complexes, namely the $\kappa$-HGLET and $\kappa$-GHWT dictionaries. The proposed MHSTs/MHSNs support multiple pooling options for the scattering transform outputs—no-pooling; local-pooling; and global-pooling—allowing them to address a wide range of problems, from classification of signals defined on simplicial complexes, to domain classification (i.e., classification of the simplicial complexes themselves), to regression of potential energies in molecular dynamics. We demonstrated that MHSNs achieve performance comparable to SotA GNNs while reducing the number of learnable parameters by up to two orders of magnitude. We attribute this success to the structure and organization of our multiscale basis dictionaries, which are naturally arranged by scales and locations, making them particularly well suited for generating scattering transform coefficients.
Future work will focus on interpreting MHST coefficients identified as important by classification methods such as the LRMs. Due to the nonlinearities in MHSTs, mapping these coefficients back to interpretable features in the primal (original) domain is challenging. To address this, we plan to investigate the optimization method proposed recently in [54], which synthesizes input patterns that maximize a model’s predicted probability for a target class, thereby revealing the meaning of selected scattering coefficients. In a related direction, we will explore how MHST coefficients can be used to identify important relationships within graphs, with the goal of narrowing the training space for attention mechanisms and graph transformers in large-scale problems.
@@ Acknowledgments
This research was partially supported by the U.S. National Science Foundation grants DMS-1912747 and CCF-1934568, and by the U.S. Office of Naval Research grant N00014-20-1-2381. N. S. also thanks Stéphane Mallat and Gabriel Peyré for their hospitality and support during his sabbatical stay (Oct. 2021–Jan. 2022) at the Centre Sciences des Données, École Normale Supérieure (Ulm); and Joan Bruna and Gabriel Peyré for inviting him to the workshop “A Multiscale tour of Harmonic Analysis and Machine Learning To Celebrate Stéphane Mallat’s 60th birthday,” (Apr. 2023) at the Institut des Hautes Études Scientifiques (IHES), Bures-sur-Yvette, where a preliminary version of this work was presented.
Appendix A A Hierarchical Bipartition Tree of a Graph
In this appendix, we summarize our construction of a hierarchical bipartition tree of a given graph $G=(V,E)$ with $|V|=n$. We will describe it for the $C_{0}$ case, but it is quite straightforward to handle the $C_{\kappa}$ ($\kappa\geq 1$) case. We impose the following four requirements for a hierarchical bipartition tree:
  1. i.
    The coarsest level is the entire graph; that is, $G_{0}^{0}=G$.
  2. ii.
    At the finest level, each region is a single node; that is, $n^{p_{\mathrm{max}}}_{k}=1$ for $0\leq k<K^{p_{\mathrm{max}}}=n$.
  3. iii.
    All regions on a given level are disjoint; that is, $V^{p}_{k}\cap V^{p}_{\tilde{k}}=\emptyset$ if $k\neq\tilde{k}$.
  4. iv.
    Each region on level $p<{p_{\mathrm{max}}}$ containing two or more nodes is partitioned into exactly two regions on level $p+1$.
One method for generating a suitable recursive partitioning of a graph is to repeatedly partition the graph and subgraphs according to the signs of their respective Fiedler vectors of $L_{\mathrm{rw}}$ matrices. However, any other method to bipartition a given graph can be used.
Appendix B Brief Description of the HGLET and the GHWT
Using a recursive partitioning of the graph, the Hierarchical Graph Laplacian Eigen Transform (HGLET) generates an overcomplete dictionary in which the supports of the basis vectors range from a single node to the entire graph [32, 55, 9]. We use $\bm{{\phi}}^{p}_{k,l}$ to denote the HGLET basis vectors, and we use $c^{p}_{k,l}$ to denote the corresponding expansion coefficients. As in the recursive partitioning, let $p\in[0,{p_{\mathrm{max}}}]$ and $k\in[0,K^{p})$ denote, respectively, the level and region to which a basis vector/coefficient corresponds. Let $l\in[0,n^{p}_{k})$ index the vectors/coefficients associated with $G^{p}_{k}$ The basis vectors are obtained by computing Laplacian eigenvectors on each subgraph $G^{p}_{k}$ and extending them by zeros to the entire graph. These may be the extended eigenvectors of $L$, $L_{\mathrm{rw}}$, or $L_{\mathrm{sym}}$ on $G^{p}_{k}$. One advantage of considering all three dictionaries is the ability to construct a hybrid basis. For example, the basis vectors for $G^{p}_{k}$ may be taken as the eigenvectors of $L_{\mathrm{sym}}(G^{p}_{k})$, whereas those for $G^{p^{\prime}}_{k^{\prime}}$ may be drawn from $L_{\mathrm{rw}}(G^{p^{\prime}}_{k^{\prime}})$. In [55], we have demonstrated the use of hybrid bases for simultaneous segmentation, denoising, and compression of classical one-dimensional signals. The computational complexity of the HGLET is $O(n^{3})$, arising from the computation of the full set of eigenvectors of the $n\times n$ Laplacian matrix at level $p=0$. When such a cost is prohibitively high, the HGLET may instead be constructed only on subgraphs $\{G^{p}_{k}\}$ with $n^{p}_{k}\leq{n_{\mathrm{max}}}<n$ nodes where, $n_{\mathrm{max}}$} is a user-specified parameter reflecting the computational budget. In this case, the complexity is reduced to $O({n_{\mathrm{max}}}^{2}n)$.
Like the HGLET, the Generalized Haar–Walsh Transform (GHWT) also employs a recursive partitioning of the graph to generate an overcomplete dictionary, but in this case the basis vectors are piecewise constant on their supports [11, 55, 9]. We denote denote the GHWT basis vectors and expansion coefficients by $\bm{{\psi}}^{p}_{k,l}$ and $d^{p}_{k,l}$, respectively. As with the HGLET, $p\in[0,{p_{\mathrm{max}}}]$ and $k\in[0,K^{p})$ denote level and region, respectively. In the GHWT framework, $l$ is referred to as the basis vector’s (or coefficient’s) tag, taking $n^{p}_{k}$ distinct values in the range $0,2^{{p_{\mathrm{max}}}-p})$. Coefficients with tag $l=0$ are called scaling coefficients, those with $l=1$ are Haar coefficients, and those with $l\geq 2$ are Walsh coefficients. For a hierarchical tree with $O(\log n)$ levels, the computational complexity of the GHWT is $O(n\log n)$.
A key feature of the GHWT is that its coefficients can be arranged in two ways. At each level $p$, grouping by their index $k$ yields the coarse-to-fine (C2F) dictionary, which has the same structure as the HGLET dictionary. Alternatively, grouping by the tag $l$ produces the fine-to-coarse (F2C) dictionary, which admits ONBs not obtainable from the C2F dictionary—for example, the true Haar ONB can be only formed only in the F2C dictionary via a graph-adapted version of the best-basis search algorithm [56]. See also [10, Sect. 6.2] and the accompanying figures for visual comparisons between the C2F and F2C GHWT dictionaries. For a graph with $n$ nodes the HGLET, GHWT-C2F, and GHWT-F2C dictionaries each contain more than $2^{\lfloor n/2\rfloor}$ choosable bases (See Table 6.1 in [57]); exceptions may occur when the recursive partitioning is highly imbalanced.
We now briefly describe theoretical properties of these dictionaries that are advantageous for signal approximation. These will be stated not in the original form given in [58, 9] for node signals, but in the updated form for signals on the $\kappa$-simplices $C_{\kappa}$ for $\kappa\in{\mathbb{Z}_{\geq 0}}$, as described in [10].
First, for singleton $\kappa$-elements $\sigma$ and $\tau$ of $C_{\kappa}$, and signal $\bm{f}$, we define a distance function and then the associated Hölder semi-norm as follows:
$d(\sigma,\tau)\,:=\,\min\left\{n^{p}_{k}\,\Big|\,\sigma,\tau\in C^{p}_{k}\right\},\quad C_{H}(\bm{f})\,:=\,\sup_{\sigma\neq\tau}\frac{\left|[\bm{f}]_{\sigma}-[\bm{f}]_{\tau}\right|}{d(\sigma,\tau)^{\alpha}}$
where $\alpha\in(0,1]$ is a fixed constant. Here, $d(\sigma,\tau)$ is the number of elements in the finest partition in the tree that constrains both $\sigma$ and $\tau$. With these definitions, the dictionary coefficient decay and approximation results of [58, 9] for the GHWT and HGLET extend directly to the $\kappa$-GHWT and $\kappa$-HGLET bases, as detailed in [10].
Theorem 3 (Fast Decay of Dictionary Coefficients as Scale Decreases).
For a simplicial complex $C$ equipped with a hierarchical bipartition tree, suppose that a signal $\bm{f}$ is Hölder continuous with exponent $\alpha\in(0,1]$ and constant $C_{H}(\bm{f})$. Then the coefficients with $l\geq 1$ for the $\kappa$-HGLET ($c^{p}_{k,l}$) and $\kappa$-GHWT ($d^{p}_{k,l}$) satisfy:
$\left|c^{p}_{k,l}\right|\leq C_{H}(\bm{f})\left(n^{p}_{k}\right)^{\alpha+\frac{1}{2}},\quad\left|d^{p}_{k,l}\right|\leq C_{H}(\bm{f})\left(n^{p}_{k}\right)^{\alpha+\frac{1}{2}}.$
Proof.
See Theorem 3.1 of [9]. ∎
Remark 2.
This theorem implies that
$\left|c^{p}_{k,l}\right|,\left|d^{p}_{k,l}\right|\leq C_{H}(\bm{f})\left(n^{p}_{k}\right)^{\alpha+\frac{1}{2}}\lessapprox C_{H}(\bm{f})\left(n/2^{p}\right)^{\alpha+\frac{1}{2}}\propto 2^{-p(\alpha+\frac{1}{2})}\propto 2^{j(\alpha+\frac{1}{2})},$
for a typical hierarchical bipartition tree of a given simplicial complex $C_{\kappa}$. Thus, these dictionary coefficients decay exponentially as the partition parameter $p$ increases and the scale parameter $j$ decreases: the finer the scale and the larger the $\alpha\in(0,1]$, the faster the decay rate. Equivalently, “the smoother the underlying function, the more rapidly these coefficients decay at finer scales.” This behavior parallels that of the conventional wavelet transforms for the Hölder continuous functions; see, e.g., [33, Sect. 6.1.3] for details.
Once an ONB from the $\kappa$-HGLET or $\kappa$-GHWT has been selected—by any method, such as the graph-adapted version of the best-basis search algorithm [56])—the following result is useful for signal approximation in that basis.
Theorem 4 (Approximation Power of Multiscale Bases).
Let $\{\bm{{\phi}}_{l}\}^{n-1}_{l=0}$ be a fixed ONB, and let $0<\rho<2$. Then,
$\|\bm{f}-P_{m}\bm{f}\|_{2}\leq\frac{|\bm{f}|_{\rho}}{m^{\beta}}.$
where $P_{m}\bm{f}$ denotes the best nonlinear $m$-term approximation of $\bm{f}$ in the basis $\{\bm{{\phi}}_{l}\}^{n-1}_{l=0}$, $\beta=\frac{1}{\rho}-\frac{1}{2}$, and $|\bm{f}|_{\rho}\,:=\,\left(\sum_{l=0}^{n-1}|\langle\bm{f},\bm{{\phi}}_{l}\rangle|^{\rho}\right)^{\frac{1}{\rho}}$.
Proof.
See Theorem 3.2 of [9] and Theorem 6.3 of [58]. ∎
Appendix C Description of Domain Classification Datasets